OpenStudy (anonymous):

How can x + 3 = -2x + 5 be set up as a system of equations?

OpenStudy (p0sitr0n):

split at equal sign y=x+3 y=-2x+5

OpenStudy (p0sitr0n):

although theres no real point in doing it since you can solve it right away

OpenStudy (anonymous):

@P0sitr0n Thank you !!

OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. -x + 5y = 1 2x + 4y = -4 2(-5y + 1) + 4y = -4 2x + 4(-5y + 1) = -4 2x + 4(5y - 1) = -4 2(5y - 1) + 4y = -4 @P0sitr0n help ?

OpenStudy (p0sitr0n):

the last one. isolate x=5y-1, replace and voila